A mass of 60 kg slides on a table. The frictional force is 60 times the velocity and the mass is


being pushed with a force of 54 sin 2t N. Find the velocity as a function of time if v = 0 at


t = 0. 


Grade:

1 Answers

Aman Bansal
12 years ago

Dear Swapnil,

F= m*a and  a= dv/dt .............

Using these two relations we can write,

accelaration of mass , = F/m - 60* integration(a)wrt time

differentiating both sides to get the expression for velocity.

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert